OpenStudy (anonymous):

A state's license plates consist of four letters followed by three numerals, and 246 letter arrangements are not allowed. How many plates can the state issue? 1 mark plates

OpenStudy (anonymous):

Please ignore the 1 mark plates at the bottom..

OpenStudy (anonymous):

total arrangements = 26^4 * 10^3 = 456976000 but since 246 aren't allowed.... we need to subtract 246*(10^3) = 246000 So answer is 456730000

OpenStudy (anonymous):

Why do you subtract 10^3??

OpenStudy (anonymous):

u subtract 246*10^3 since for every one of the 246 letter arrangements, there are 1000 different number combinations that can follow....so 246000 total combinations that are not allowed.